# Hubble Tension or Distance Ladder Crisis?

### v3

@@ Abstract
We present an up-to-date compilation of published Hubble constant ($H_{0}$) measurements that are independent of the CMB sound horizon scale. This compilation is split in two distinct groups: A. Distance Ladder Measurements sample comprising of 19 recent measurements, primarily from the past four years, utilizing various rung 2 calibrators and rung 3 cosmic distance indicators. B.One-Step Measurements sample including 31 measurements of $H_{0}$ that are independent of both the CMB sound horizon scale and the distance ladder approach. These 31 measurements are derived from diverse probes such as Cosmic Chronometers, gamma-ray attenuation, strong lensing, megamasers etc. Statistical analysis reveals a significant distinction between the two samples. The distance ladder-based sample yields a best fit $H_{0}=72.76\pm 0.50$ km s-1 Mpc-1 with $\chi^{2}/dof=0.52$ indicating some correlations. The one-step measurements result in $H_{0}=69.01\pm 0.49$ km s-1 Mpc-1 with $\chi^{2}/dof=1.43$ indicating some internal tension. If two outlier measurements are removed (TDCOSMO.I-2019 known to have systematics and MCP-2020) the best fit of the one step sample reduces to $H_{0}=68.3\pm 0.5$ km s-1 Mpc-1 with $\chi^{2}/dof=1.01$, fully self-consistent and consistent with sound horizon based measurements. A Kolmogorov-Smirnov test yields a p-value of 0.0001 suggesting that the two samples are fundamentally distinct, with a probability of less than 0.01% that they are drawn from the same underlying distribution. These findings suggest that the core of the Hubble tension lies not between early and late-time measurements, but between distance ladder measurements and all other $H_{0}$ determinations. This discrepancy points to either a systematic effect influencing all distance ladder measurements or a fundamental physics anomaly affecting at least one rung of the distance ladder.
@@ I Introduction
The main goal of the present analysis is to address these questions using an extensive up-to-date compilation of recent measurements of $H_{0}$ (most published during the past 5 years with emphasis in the last 3 years) which are independent of the sound horizon scale and include both distance ladder measurements and one step measurements that involve no rungs. We thus split the measurements of $H_{0}$ that are independent of the sound horizon scale in two groups: 1. Distance ladder rung based measurements 2. One step measurements that involve direct distance or $H(z)$ probes. We then test the statistical properties of the $H_{0}$ measurements in each group in order to address the above questions. This is the first time that such a statistical meta-analysis is attempted since previous corresponding studies have not separated the two groups of measurements and have also included many early outdated measurementsWang et al. (2023); Lopez-Corredoira (2022); Cao and Ratra (2023); Verde et al. (2023).
@@ II.1.2 Measuring $H_{0}$ from the CMB
The angular scale of the first acoustic peak in the CMB power spectrum, $\theta_{s}$, is inversely proportional to the sound horizon at the surface of last scattering:
@@ II.1.3 Connecting BAO and the CMB
Building on the CMB measurements discussed in the previous subsection, Baryon Acoustic Oscillations (BAO) provide an independent confirmation of the sound horizon scale. The BAO feature serves as a standard ruler in the late-time universe, manifesting as a peak in the correlation function of galaxies. This peak corresponds to the same comoving scale $r_{s}$ imprinted in the CMB.
@@ 3. Standardizing SN Ia luminosities using the Tripp relation Tripp (1998):
    • Examples: Galaxies hosting both Cepheids/TRGB/Miras and SnIa
@@ II.3.5 Cosmic Chronometers
    Hubble parameter estimation: Applying Equation 40 to calculate $H(z)$ at various redshifts.
@@ II.3.6 Standard Sirens and Kilonovae
Standard sirens, a term coined for gravitational waves (GWs) emitted from compact binary mergers, provide a unique method for measuring the Hubble constant, $H_{0}$, directly from the gravitational wave signal without requiring a cosmic distance ladder Abbott et al. (2017); Hotokezaka et al. (2019); Chen et al. (2018); Ando et al. (2013); Bulla et al. (2022). This section details the principles behind using GWs and their electromagnetic counterparts, kilonovae, as cosmological probes, highlighting their potential to resolve the Hubble tension.
@@
Gravitational waves emitted from mergers of binary neutron stars (BNS) or neutron star-black hole (NS-BH) binaries offer a direct measure of the luminosity distance to the source. The amplitude of the gravitational wave signal is inversely proportional to this distance, and by combining it with the redshift of the host galaxy, $H_{0}$ can be directly inferred.
@@
The landmark event GW170817, where both gravitational waves and a kilonova (AT 2017gfo) were observed, serves as a prime example of this method. The gravitational wave signal provided a measurement of the luminosity distance, while the identification of the host galaxy NGC 4993 led to a redshift measurement. Combining these measurements allowed for an independent estimate of $H_{0}$.
@@ 1 Reese (2003) Reese (2003) $61\pm 18$ SZ Effect
24-25. Pogosian (2024) Pogosian et al. (2024): This analysis presents a cosmological-model-independent measurement of the Hubble constant $H_{0}$ using the DESI Year 1 BAO data combined with the CMB acoustic scale $\theta_{*}$ and a Planck $\Lambda$CDM prior on $\Omega_{m}h^{2}$. The authors treat the sound horizon at baryon decoupling $r_{d}$ as a free parameter, avoiding assumptions about recombination physics. Using CosmoMC, they find $H_{0}=69.88\pm 0.93$ km s-1 Mpc-1, which is $\sim 2\sigma$ higher than the Planck $\Lambda$CDM value. For comparison, they perform the same analysis with pre-DESI BAO data, obtaining $H_{0}=67.37\pm 0.96$ km s-1 Mpc-1. The difference stems from the larger $r_{d}h$ value measured by DESI. This method provides a valuable consistency test of the cosmological model at recombination, independent of specific recombination physics assumptions. The authors critically discuss the differences between DESI and pre-DESI results, noting that future DESI data covering larger volumes will help resolve discrepancies. While this approach does not fully resolve the Hubble tension, it offers important insights into potential modifications needed in the standard cosmological model around the epoch of recombination.
@@ III.2 Results: One step methods vs distance ladder methods
    • Distance Ladder Dependent: $H_{0}=72.76\pm 0.50$ km/s/Mpc
    • One-step (Distance Ladder Independent): $H_{0}=69.01\pm 0.49$ km/s/Mpc. This value decreases further to $H_{0}=68.3\pm 0.5$ km/s/Mpc if two outliers are removed from the sample as discussed below.
@@ III.2.1 $\chi^{2}$ analysis and robustness of results
Table 3: Analysis of Various Subsets of the One Step - Sound Horizon and Distance Ladder Independent Sample. The data numbers correspond to Table 2. Notice the dramatic improvement of the self consistency of the full sample (decrease of the reduced $\chi^{2}_{\text{red}}\equiv\chi^{2}/dof$ ) if the two outliers (TDCOSM.I and MCP-SH0ES) are removed.
@@
Full Sample $69.0\pm 0.5$ 43.14 1.44 30 All
Full Sample (excl. 3,4) $68.3\pm 0.5$ 28.16 1.01 28 All except 3,4
GW and Kilonovae EPM $66.3\pm 2.0$ 8.13 1.16 7 6,9,14,15,17,18,22,27
Lensing All $71.4\pm 1.0$ 12.00 1.33 9 3,5,11,12,19,20,26,29-31
Lensing (no TDCOSMO I) $69.7\pm 1.2$ 7.14 0.89 8 5,11,12,19,20,26,29-31
Cosmic Chronometers $68.3\pm 2.0$ 1.80 0.90 2 8,10,13
Megamasers $72.3\pm 2.7$ 1.39 1.39 1 2,4
@@ IV Conclusion-Discussion-Future Prospects
This study presents a comprehensive analysis of Hubble constant ($H_{0}$) measurements, categorizing them into two distinct groups: distance ladder measurements and one-step measurements. The distance ladder approach, which relies on various calibrators and cosmic distance indicators, yielded an $H_{0}$ value of $72.76\pm 0.50$ km/s/Mpc. In contrast, one-step measurements, which are independent of the CMB sound horizon scale and do not require a distance ladder, indicated a lower $H_{0}$ of $69.01\pm 0.49$ km/s/Mpc. It’s important to note that these uncertainties are likely underestimated, as the $H_{0}$ measurements were assumed to be independent, an assumption not fully realized in practice.
@@
    • Modified gravity theories with scale-dependent effects Perivolaropoulos and Skara (2022c); Desmond and Sakstein (2020); Desmond et al. (2019).

### v4

@@ Abstract
We present an up-to-date compilation of published Hubble constant ($H_{0}$) measurements that are independent of the CMB sound horizon scale. This compilation is split in two distinct groups: A. Distance Ladder Measurements sample comprising of 20 recent measurements, primarily from the past four years, utilizing various rung 2 calibrators and rung 3 cosmic distance indicators. B.One-Step Measurements sample including 33 measurements of $H_{0}$ that are independent of both the CMB sound horizon scale and the distance ladder approach. These 33 measurements are derived from diverse probes such as Cosmic Chronometers, gamma-ray attenuation, strong lensing, megamasers etc. Statistical analysis reveals a significant distinction between the two samples. The distance ladder-based sample yields a best fit $H_{0}=72.8\pm 0.5$ km s-1 Mpc-1 with $\chi^{2}/dof=0.51$ indicating some correlations. The one-step measurements result in $H_{0}=69.0\pm 0.48$ km s-1 Mpc-1 with $\chi^{2}/dof=1.37$ indicating some internal tension. If two outlier measurements are removed (TDCOSMO.I-2019 known to have systematics and MCP-2020) the best fit of the one step sample reduces to $H_{0}=68.3\pm 0.5$ km s-1 Mpc-1 with $\chi^{2}/dof=0.95$, fully self-consistent and consistent with sound horizon based measurements. A Kolmogorov-Smirnov test yields a p-value of 0.0001 suggesting that the two samples are fundamentally distinct, with a probability of less than 0.01% that they are drawn from the same underlying distribution. These findings suggest that the core of the Hubble tension lies not between early and late-time measurements, but between distance ladder measurements and all other $H_{0}$ determinations. This discrepancy points to either a systematic effect influencing all distance ladder measurements or a fundamental physics anomaly affecting at least one rung of the distance ladder.
@@ I Introduction
The main goal of the present analysis is to address these questions using an extensive up-to-date compilation of recent measurements of $H_{0}$ (most published during the past 5 years with emphasis in the last 3 years) which are independent of the sound horizon scale and include both distance ladder measurements and one step measurements that involve no rungs. We thus split the measurements of $H_{0}$ that are independent of the sound horizon scale in two groups: 1. Distance ladder rung based measurements 2. One step measurements that involve direct distance or $H(z)$ probes and are independent of the recombination sound horizon scale. We then test the statistical properties of the $H_{0}$ measurements in each group in order to address the above questions. This is the first time that such a statistical meta-analysis is attempted since previous corresponding studies have not separated the two groups of measurements and have also included many early outdated measurementsWang et al. (2023); Lopez-Corredoira (2022); Cao and Ratra (2023); Verde et al. (2023).
@@ II.1.2 Measuring $H_{0}$ from the CMB
The angular scale of the first acoustic peak in the CMB power spectrum, $\theta_{s}$, is inversely proportional to the sound horizon at the surface of last scatteringEfstathiou (2001); Spergel et al. (2003); Aghanim et al. (2020); Choi et al. (2020):
@@ II.1.3 Connecting BAO and the CMB
Building on the CMB measurements discussed in the previous subsection, Baryon Acoustic Oscillations (BAO) provide an independent confirmation of the sound horizon scaleEisenstein et al. (2005); Aubourg et al. (2015); Bautista et al. (2020); Favale et al. (2024). The BAO feature serves as a standard ruler in the late-time universe, manifesting as a peak in the correlation function of galaxies. This peak corresponds to the same comoving scale $r_{s}$ imprinted in the CMB.
@@ 3. Standardizing SN Ia luminosities using the Tripp relation Tripp (1998):
    • Examples: Galaxies hosting both Cepheids/TRGB/Miras/JAGB and SnIa
@@ II.3.5 Cosmic Chronometers
    Hubble parameter estimation: Applying Equation (40) to calculate $H(z)$ at various redshifts.
@@ II.3.6 Standard Sirens and Kilonovae
Standard sirens, a term coined for gravitational waves (GWs) emitted from compact binary mergers, provide a unique method for measuring the Hubble constant, $H_{0}$, directly from the gravitational wave signal without requiring a cosmic distance ladder Abbott et al. (2017a, b); Hotokezaka et al. (2019); Chen et al. (2018); Ando et al. (2013); Bulla et al. (2022). This section details the principles behind using GWs and their electromagnetic counterparts, kilonovae, as cosmological probes, highlighting their potential to resolve the Hubble tension.
@@
Gravitational waves emitted from mergers of binary neutron stars (BNS) or neutron star-black hole (NS-BH) binaries offer a direct measure of the luminosity distance to the sourceAbbott et al. (2017a). The amplitude of the gravitational wave signal is inversely proportional to this distance, and by combining it with the redshift of the host galaxy, $H_{0}$ can be directly inferred.
@@
The landmark event GW170817, where both gravitational waves and a kilonova Abbott et al. (2017c, a) were observed, serves as a prime example of this method. The gravitational wave signal provided a measurement of the luminosity distance, while the identification of the host galaxy NGC 4993 led to a redshift measurement. Combining these measurements allowed for an independent estimate of $H_{0}$.
@@ III.1.1 Distance Ladder-Based Measurements
19 Boubel (2024) Boubel et al. (2024) $73.3\pm 4.08$ Tully Fisher + Cepheid + TRGB
20 Said (2024) Said et al. (2024) $76.05\pm 4.90$ DESI FP relation + SBF
@@
20. Said et al. (2024) Said et al. (2024): This analysis employed the Fundamental Plane (FP) relation for early-type galaxies in combination with Surface Brightness Fluctuation (SBF) measurements to estimate the Hubble constant. The study utilized data from the Dark Energy Spectroscopic Instrument (DESI) Peculiar Velocity Survey, focusing on a sample of early-type galaxies. The FP relation, which correlates the effective radius, velocity dispersion, and surface brightness of elliptical galaxies, was calibrated using SBF distances to nearby galaxies, particularly the Coma cluster. This approach yielded $H_{0}=76.05\pm 4.90$ km s-1 Mpc-1. It’s important to note that this method is dependent on the cosmic distance ladder and requires local calibration. The reliance on local calibrators, such as SBF measurements of the Coma cluster, means that the resulting $H_{0}$ value is sensitive to potential systematic errors in these calibrations. This dependence on local distance indicators places this measurement in the category of distance ladder dependent $H_{0}$ determinations, which often yield higher values compared to measurements that are independent of the distance ladder.
@@ 1 Reese (2003) Reese (2003) $61\pm 18$ SZ Effect
2. Kuo 2013 Kuo et al. (2013): This analysis utilized the water megamaser technique to measure the angular diameter distance to NGC 6264, a galaxy located at approximately 140 Mpc. This study is notable for being one of the first to apply the megamaser method to a galaxy well into the Hubble flow, where peculiar velocities have a reduced impact on distance measurements. The authors used Very Long Baseline Interferometry (VLBI) observations to map the spatial and velocity distribution of masers in the circumnuclear disk, combined with single-dish monitoring to measure accelerations. By modeling the three-dimensional structure of the maser disk, they derived a Hubble constant of $H_{0}=68\pm 9$ km s-1 Mpc-1. This result is in tension with a later study by Pesce et al. (2020), which found $H_{0}=73.9\pm 3.0$ km s-1 Mpc-1 using a larger sample of megamasers. However, the Pesce et al. sample included galaxies at smaller distances, ranging from about 8.5 to 140 Mpc ($z\approx 0.002-0.034$), potentially making it more susceptible to biases from peculiar velocities. While the Kuo et al. measurement has larger uncertainties due to the challenges of observing more distant masers, it benefits from reduced systematic errors associated with local flows, making it an important independent check on $H_{0}$ measurements derived from nearer galaxies.
@@
8. Wu et al. (2021)Wu et al. (2022): This analysis employed a novel method using Fast Radio Bursts (FRBs) to measure the Hubble constant, independent of the cosmic distance ladder. This approach leverages the dispersion measure (DM) of FRBs, which is related to the electron column density along the line of sight and thus serves as a proxy for cosmological distance. The authors used a sample of 18 localized FRBs with known redshifts and carefully modeled the contributions to the DM from the Milky Way, host galaxies, and the intergalactic medium (IGM). By using the DM-redshift relation and accounting for the probability distributions of various DM components, they derived $H_{0}=64.67^{+5.62}_{-4.66}$ km s-1 Mpc-1. This method is particularly noteworthy as it requires no local distance calibrators and is independent of the cosmic distance ladder, potentially offering a new probe to address the Hubble tension. The authors also demonstrated that with a future sample of 100 localized FRBs, the precision of $H_{0}$ measurement could reach $\sim$2.6%, making FRBs a promising tool for cosmology.
@@
26-27. Pogosian (2024) Pogosian et al. (2024): This analysis presents a cosmological-model-independent measurement of the Hubble constant $H_{0}$ using the DESI Year 1 BAO data combined with the CMB acoustic scale $\theta_{*}$ and a Planck $\Lambda$CDM prior on $\Omega_{m}h^{2}$. The authors treat the sound horizon at baryon decoupling $r_{d}$ as a free parameter, avoiding assumptions about recombination physics. Using CosmoMC, they find $H_{0}=69.88\pm 0.93$ km s-1 Mpc-1, which is $\sim 2\sigma$ higher than the Planck $\Lambda$CDM value. For comparison, they perform the same analysis with pre-DESI BAO data, obtaining $H_{0}=67.37\pm 0.96$ km s-1 Mpc-1. The difference stems from the larger $r_{d}h$ value measured by DESI. This method provides a valuable consistency test of the cosmological model at recombination, with no direct dependence on the sound horizon scale. It should be stressed however that the assumed prior on $\Omega_{m}h^{2}$ is model dependent assumes $\Lambda$CDM and would be violated in the context of EDE models which predict a higher value of $\Omega_{m}h^{2}$oulin et al. (2023). The authors critically discuss the differences between DESI and pre-DESI results, noting that future DESI data covering larger volumes will help resolve discrepancies. While this approach does not fully resolve the Hubble tension, it offers important insights into potential modifications needed in the standard cosmological model around the epoch of recombination.
@@ III.2 Results: One step methods vs distance ladder methods
    • Distance Ladder Dependent: $H_{0}=72.8\pm 0.5$ km/s/Mpc
    • One-step (Distance Ladder Independent): $H_{0}=69.0\pm 0.5$ km/s/Mpc. This value decreases further to $H_{0}=68.3\pm 0.5$ km/s/Mpc if two outliers are removed from the sample as discussed below.
@@ III.2.1 $\chi^{2}$ analysis and robustness of results
Table 3: Analysis of Various Subsets of the One Step - Sound Horizon and Distance Ladder Independent Sample. The data numbers correspond to Table 2. Notice the dramatic improvement of the self consistency of the full sample (decrease of the reduced $\chi^{2}_{\text{red}}\equiv\chi^{2}/dof$ ) if the two outliers (TDCOSMO.I and MCP-SH0ES) are removed. These outliers have been removed from most of the subsamples after row 2. If they were present the value of $\chi^{2}_{red}$ would have been significantly larger (see eg row 5 where TDCOSMO.I has been included).
@@
Full Sample $69.0\pm 0.5$ 43.69 1.37 32 All
Full Sample (excl. 4,5) $68.3\pm 0.5$ 28.51 0.95 30 All except outliers 4,5
Gravitational Waves (GW) $71.4\pm 3.1$ 1.18 0.24 5 7,11,16,17,24,29
Full Sample (excl. GW) $68.2\pm 0.5$ 26.31 1.10 24 All except 4,5,7,11,16,17,24,29
TD Lensing All $71.4\pm 1.0$ 12.00 1.33 9 4,6,13,14,21,22,28,31-33
Full Sample (excl. TD Lensing) $68.2\pm 0.6$ 23.46 1.07 22 All except 4-6,13,14,21,22,28,31-33
TD Lensing (no TDCOSMO I) $69.7\pm 1.2$ 7.14 0.89 8 6,13,14,21,22,28,31-33
Cosmic Chronometers $68.3\pm 2.0$ 1.80 0.90 2 10,12,15
Full Sample (excl. CC) $68.3\pm 0.5$ 26.71 0.99 27 All except 4,5,10,12,15
Megamasers $72.0\pm 2.6$ 1.60 0.80 2 2,3,5
Full Sample (excl. Megamasers) $68.3\pm 0.5$ 28.36 1.01 28 All except 2-5
Megamasers (excl. MCP-SH0ES) $67.0\pm 5.0$ 0.03 0.03 1 2,3
Early Time (No Sound Horizon) $68.4\pm 0.6$ 5.93 2.97 2 9,26,27
Full Sample (excl. Early Time) $68.1\pm 0.9$ 22.50 0.83 27 All except 4,5,9,26,27
@@ IV Conclusion-Discussion-Future Prospects
This study presents a comprehensive analysis of Hubble constant ($H_{0}$) measurements, categorizing them into two distinct groups: distance ladder measurements and one-step measurements. The distance ladder approach, which relies on various calibrators and cosmic distance indicators, yielded an $H_{0}$ value of $72.8\pm 0.5$ km/s/Mpc. In contrast, one-step measurements, which are independent of the CMB sound horizon scale and do not require a distance ladder, indicated a lower $H_{0}$ of $69.0\pm 0.5$ km/s/Mpc. It’s important to note that these uncertainties are likely underestimated, as the $H_{0}$ measurements were assumed to be independent, an assumption not fully realized in practice.
@@
    • Modified gravity theories with scale-dependent effects Perivolaropoulos and Skara (2022c); Desmond and Sakstein (2020); Desmond et al. (2019); Högås and Mörtsell (2023).